a rectangle has it breath to be 2x and x+y-11.but it lenght is 3y+2 and x+7. find the value of x and y and find the dimesion of the rectangle
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

?

Makes no sense.
(Both dimensions negative)

Answer by ikleyn(52781)

.
From the condition, we have these two equations


2x = x + y -11    (1)
3y + 2 = x+ 7     (2)


Simplify each equation


 x -  y = -11      (1')
-x + 3y =   5      (2')


Add equations (1') and (2'). You will get

    2y = -11 + 5 = -6  ====>  y = -6/2 = -3.


Then from eq(1')  x = y - 11 = -3 - 11 = -14.


Answer.  x= -14;  y= -3.

Solved.

-----------------

@josgarithmetic correctly noticed that with these values of x and y the dimensions of the "rectangle" are negative,

so, such a "rectangle" does not exist.

It is true.


=====================

Every night we have a visitor in the forum who posts frank/candid RUBBISH.